2T^-4 -5T^-2 +2=0



2T HAS A NEGATIVE EXPONENT OF -4 AND 5T HAS A NEGATIVE EXPONENT OF -2

solve





thanks

Answer by stanbon(75887)

2T^-4 -5T^-2 +2=0
--------------
Let x = T^-2
You get:
2x^2 - 5x + 2 = 0
(2x-1)(x-2)=0
x = 1/2 or x = 2
----------------
Therefore:
T^-2 = 1/2 or T^-2 = 2
T = (1/2)^(-1/2) or T = 2^(-1/2)
T = sqrt2 or T = (sqrt2)/2
=========================
Cheers,
Stan H.